CashMoneyDay:

Factor the expression x2+2x−15

CripQUEZZ:

Can you post a ss plz

CashMoneyDay:

1 attachment
CripQUEZZ:

I’d say it’s B bc X + 3 = 3x And then X + 12 = 15 So I’m guessing that x = 3 So it would be 3x^2 + 15

CashMoneyDay:

what are the steps to get ther though

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ I’d say it’s B bc X + 3 = 3x And then X + 12 = 15 So I’m guessing that x = 3 So it would be 3x^2 + 15 \(\color{#0cbb34}{\text{End of Quote}}\) And also x^2 + 2x = 3x^2

XioGonz:

When Factoring x^2+2x-15 First term is: x2 coefficient is 1. Middle term: +2x coefficient is 2 . Last term:, the constant is -15 Multiply the coefficient of the first term by the constant 1 • -15 = -15 Find two factors of -15 which has the sum equal to the coefficient of the middle term (Again it is 2) -15+1=-14 -5+3=-2 -3+5=2 x^2 - 3x + 5x - 15 x • (x-3) 5 • (x-3) Hence, (x+5) • (x-3)

CashMoneyDay:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz When Factoring x^2+2x-15 ohhhh okay First term is: x2 coefficient is 1. Middle term: +2x coefficient is 2 . Last term:, the constant is -15 Multiply the coefficient of the first term by the constant 1 • -15 = -15 Find two factors of -15 which has the sum equal to the coefficient of the middle term (Again it is 2) -15+1=-14 -5+3=-2 -3+5=2 x^2 - 3x + 5x - 15 x • (x-3) 5 • (x-3) Hence, (x+5) • (x-3) \(\color{#0cbb34}{\text{End of Quote}}\)

XioGonz:

...

SemiDefinite:

learn to see patterns, or give in to the Bots